# Set Features for Anomaly Detection

### v3

@@ Abstract
This paper proposes to use set features for detecting anomalies in samples that consist of unusual combinations of normal elements. Many leading methods discover anomalies by detecting an unusual part of a sample. For example, state-of-the-art segmentation-based approaches, first classify each element of the sample (e.g., image patch) as normal or anomalous and then classify the entire sample as anomalous if it contains anomalous elements. However, such approaches do not extend well to scenarios where the anomalies are expressed by an unusual combination of normal elements. In this paper, we overcome this limitation by proposing set features that model each sample by the distribution of its elements. We compute the anomaly score of each sample using a simple density estimation method, using fixed features. Our approach [^1] The code is available on github: https://github.com/NivC/SINBAD outperforms the previous state-of-the-art in image-level logical anomaly detection and sequence-level time series anomaly detection.
@@ 1 Introduction
Anomaly detection aims to automatically identify samples that exhibit unexpected behavior. In some anomaly detection tasks anomalies are quite subtle. For example, let us consider an image of a bag containing screws, nuts, and washers (Fig.1). There are two ways in which a sample can be anomalous: (i) one or more of the elements in the sample are anomalous. E.g., a broken screw. (ii) the elements are normal but appear in an anomalous combination. E.g., one of the washers might be replaced with a nut.
In recent years, remarkable progress has been made in detecting samples featuring anomalous elements. Segmentation-based methods were able to achieve very strong results on industrial inspection datasets [3]. Such methods operate in two stages: First, we perform anomaly segmentation by detecting which (if any) of the elements of the sample are anomalous, e.g., by density estimation [13, 17, 48]. Given an anomaly segmentation map, we compute the sample-wise anomaly score as the number of anomalous elements, or the abnormality level of the most anomalous element. If the anomaly score exceeds a threshold, the entire sample is denoted as an anomaly. We denote this paradigm detection-by-segmentation.
@@
We propose to detect anomalies consisting of unusual combinations of normal elements using set representations. The key insight in this work, that we should treat a sample as the set of its elements, is driven by the assumption that in many cases the distribution of elements in a sample is more correlated with it being anomalous than with the ordering of the elements. Each sample is therefore modeled as an orderless set. The elements are represented using standard fixed feature embeddings, e.g., a deep representation extracted by a pre-trained neural network or handcrafted features. To describe this set of features we model their distrbution as a set using a collection of histograms. We compute a histogram for a collection of random projection directions in feature space. The bin occupancies from all the histograms are concatenated together, forming our set representation. Finally, we score anomalies using density estimation on this set representation. We compare our set descriptor to previous approaches and highlight its connection to the sliced Wasserstein distance (SWD).
Our method, SINBAD (Set INspection Based Aomalies Detection) is evaluated on two diverse tasks. The first task is image-level logical anomaly detection on the MVTec-LOCO datasets. Our method outperforms more complex state-of-the-art methods, while not requiring any training. We also evaluate our method on series-level time series anomaly detection. Our approach outperforms current methods on logical anomalies while not using augmentations or training. Note that our method relies on the prior that the elements are normal but their combination is anomalous. In scenarios where the elements themselves are anomalous, it is typically better to perform anomaly detection directly at the element level (e.g., detection-by-segmentation or other methods).
@@ 2 Previous work
Image Anomaly Detection. A comprehensive review of anomaly detection can be found in [52]. Early approaches ([21, 34, 20]) used handcrafted representations. Deep learning has provided a significant improvement on such benchmarks [33, 50, 22, 27, 51, 40, 53, 59]. As density estimation methods utilizing pre-trained deep representation have made significant steps towards the supervised performance on such benchmarks [16, 12, 45, 44, 46], much research is now directed at other challenges [46]. Such challenges include detecting anomalous image parts which are small and fine-grained [13, 36, 17, 48, 28]. The progress in anomaly detection and segmentation has been enabled by the introduction of appropriate datasets [3, 4, 11, 30, 7]. Recently, the MVTec-LOCO dataset [5] has put the spotlight on fine-grained anomalies that cannot be identified using single patches, but can only be identified when examining the connection between different (otherwise normal) elements in an image. Here, we will focus on detecting such logical anomalies.
@@
Time series Anomaly detection. A general review on anomaly detection in time series can be found in [6]. In this paper, we are concerned with anomaly detection of entire sequences, i.e., cases where an entire signal may be abnormal. Traditional approaches for this task include generic anomaly detection approaches such as $k$ nearest neighbors ($k$NN) based methods, e.g., vanilla $k$NN [20] and Local Outlier Factor (LOF) [10], Tree-based methods [37], One-class classification methods [60] and SVDD [56], and auto-regressive methods that are particular to time series anomaly detection [49]. With the advent of deep learning, the traditional approaches were augmented with deep-learned features: Deep one-class classification methods include DeepSVDD [50] and DROCC [24]. Deep auto-regressive methods include RNN-based prediction and auto-encoding methods [9, 38]. In addition, some deep learning anomaly detection approaches are conceptually different from traditional approaches. These methods use classifiers trained on normal data, assuming they will struggle to generalize to anomalous data [2, 42].
@@
Discretized Projections. Discretized projections of multivariate data have been used in many previous works. Locally sensitive hashing [15] uses random projection and subsequent binary quantization as a hash for high-dimensional data. It was used to facilitate fast $k$ nearest neighbor search. Random projections transformation is also highly related to the Radon transform [43]. Kolouri et al. [32] used this representation as a building block in their set representation. HBOS [23] performs anomaly detection by representing each dimension of multivariate data using a histogram of discretized variables. LODA [41] extends this work, by first projecting the data using a random projection matrix. We differ from LODA in the use of a different density estimator and in using sets of multiple elements rather than single sample descriptions. Rocket and mini-rocket [18, 19] represent time series for classification using the averages of their window projection.
@@ 3.1 A Set is More Than the Sum of Its Parts
Detecting anomalies in complex samples consisting of collections of elements requires understanding how the different elements of each sample interact with one another. As a motivating example let us consider the screw bag class from the MVTec-LOCO dataset (Fig. 1). Each normal sample in this class contains two screws (of different lengths), two nuts, and two washers. Anomalies may occur, for example, when an additional nut replaces one of the washers. Detecting anomalies such as these requires describing of all elements within a sample together, since each local element on its own could have come from a normal sample.
A typical way to aggregate element descriptor features is by average pooling. Yet, this is not always suitable for set anomaly detection. In supervised learning, average pooling is often built into architectures such as ResNet [25] or DeepSets [61], in order to aggregate local features. Therefore, deep features learnt with a supervised loss are already trained to be effective for pooling. However, for lower-level feature descriptors this may not be the case. As demonstrated in Fig.2. The average of a set of features is far from a complete description of the set. This is especially true in anomaly detection, where density estimation approaches require more discriminative features than those needed for supervised learning [46]. Even when an average pooled set of features work for a supervised task, it might not work for anomaly detection.
Therefore, we choose to model a set by the distribution of its elements, ignoring the ordering between them. A naive way of doing so is using a discretized, volumetric representation, similarly to $3$D voxels for point clouds. Unfortunately, such approaches cannot scale to high dimensions, and more compact representations are required. We choose to represent sets using a collection of $1$D histograms. Each histogram represents the density of the elements of the set when projected along a particular direction. We provide an illustration of this idea in Figure 2.
In some cases, projecting a set along its original axes may not be discriminative enough. Histograms along the original axes correspond to 1D marginals, and may map distant elements to the same histogram bins (see Fig.2 for an illustration). On the other side, we can see at the bottom of the figure that when the set elements are first projected along another direction, the histograms of the two sets are distinct. This suggests a set description method: first project each set along a shared random direction and then compute a 1D histogram for each set along this direction. We can obtain a more powerful descriptor by repeating this procedure with projections along multiple random directions. We analyze this approach in Section 3.5
@@ 3.2 Preliminaries
We are provided a training set $\mathcal{S}$ containing a set of $N_{S}$ samples $x_{1},x_{2}..x_{N_{S}}\in\mathcal{S}$. All the samples at training time are known to be normal. At test time, we are presented with a new sample $\tilde{x}$. Our objective is to learn a model, which operates on each sample $\tilde{x}$ and outputs an anomaly score. Samples with anomaly scores higher than a predetermined threshold value are labeled as anomalies. The unique aspect of our method is its treatment of each sample $x$ as consisting of a set of $N_{E}$ elements $x=[e_{1},e_{2}..e_{N_{E}}]$. Examples of such elements include patches for images, or a temporal windows for time series. We assume the existence of a powerful feature extractor $F$ that maps each raw element $e$ into an element feature descriptor $f_{e}$. We will describe specific implementations of the feature extraction for two important applications: images and time series, in Sec. 4.
@@ 3.3 Set Features by Histogram of Projections
Motivated by the toy example in Sec. 3.1, we propose to model each set $x$ by the histogram of the values of its elements along a collection of directions.
Histogram descriptor. Average pooling of the features of all elements in the set may result in insufficiently informative representations (Sec. 3.1). Instead, we describe the set of features using the histogram of values along a collection of directions. We note the set $s$ composed of the set of values of each sample $f$ in each dimension $j$ as $s[j]=\{f_{1}[j],f_{2}[j]..f_{N_{S}}[j]\}$. We compute the maximal and minimal values for $s[j]$ each dimension $j$, and divide the region between them into $K$ bins. We compute histograms $H_{j}$ for each of the $N_{D}$ dimensions, describing the set $s$, and concatenate them into a single set descriptor $h$. The descriptor of each set therefore has a dimension of $N_{D}\cdot K$.
Projection. As discussed before, not all projection directions are equally informative for describing the distributions of sets. In the general case, it is unknown which directions will be the most informative ones for capturing the difference between normal and anomalous sets. As we cannot tell the best projection directions in advance, we randomly project the features. This ensures a low likelihood for catastrophically poor projection directions, such as those in the example in Fig.2.
In practice, we generate a random projection matrix $P\in R^{(N_{D},N_{P})}$ by sampling values for each dimension from the Gaussian distribution $\mathbb{N}(0,1)$. We project the features $f$ of each element of $x$, yielding projected features $f^{\prime}$:
$f^{\prime}=Pf$ (1)
We run the histogram descriptor procedure described above on the projected features. The final set descriptor $h_{Px}$ becomes the concatenation of $N_{P}$ histograms, resulting in a dimension of $N_{P}\cdot K$.
@@ 3.5 Connection to Previous Set Descriptors and the Wasserstein Distance
Classical set descriptors. Many prior methods have been used to describe sets of image elements, among them Bag-of-Features [14], VLAD [29], and Fisher-Vectors [54]. These begin with a preliminary clustering stage (K-means or Gaussian Mixture Model). They then describe the set using the zeroth, first, or second moments of each cluster. The comparison in Section 12 shows that our method outperforms clustering-based methods in describing our feature sets.
Wasserstein distance. Our method is closely related to the Wasserstein distance, which measures the minimal distance required to transport the probability mass from one distribution to the other. As computing the Wasserstein distance for high-dimensional data such as ours is computationally demanding, the Sliced Wasserstein Distance (SWD) [8], was proposed as an alternative. The $SWD_{1}$ between two sets. $x$ and $y$, has a particularly simple form:
@@
As the histogram projection have high correlation between them, it is necessary to decorrelate them. This is done here using a Gaussian model. The Mahalanobis distance used here therefore performs better than the simple $SWD_{1}$ distance. While this weakens the connection to the Wasserstein distance, this was crucial for most time-series datasets (see Table 13). In practice, we opted to use $k$NN with the Mahalanobis distance rather than simply computing the Mahalanobis distance to $\mu$ as it worked slightly better (see Section 12).

### v4

@@ Abstract
This paper proposes to use set features for detecting anomalies in samples that consist of unusual combinations of normal elements. Many leading methods discover anomalies by detecting an unusual part of a sample. For example, state-of-the-art segmentation-based approaches, first classify each element of the sample (e.g., image patch) as normal or anomalous and then classify the entire sample as anomalous if it contains anomalous elements. However, such approaches do not extend well to scenarios where the anomalies are expressed by an unusual combination of normal elements. In this paper, we overcome this limitation by proposing set features that model each sample by the distribution of its elements. We compute the anomaly score of each sample using a simple density estimation method, using fixed features. Our approach outperforms the previous state-of-the-art in image-level logical anomaly detection and sequence-level time series anomaly detection [^1] Our code can be found at https://github.com/NivC/SINBAD/.
@@ 1 Introduction
Anomaly detection aims to automatically identify samples that exhibit unexpected behavior. In some detection tasks anomalies are quite subtle. For example, let us consider an image of a bag containing screws, nuts, and washers (Fig.1). There are two ways in which a sample can be anomalous: (i) one or more of the elements in the sample are anomalous. E.g., a broken screw. (ii) the elements are normal but appear in an anomalous combination. E.g., one of the washers might be replaced with a nut.
In recent years, remarkable progress has been made in detecting samples featuring anomalous elements. Segmentation-based methods were able to achieve very strong results on industrial inspection datasets (Bergmann et al., 2019). Such methods operate in two stages: First, we perform anomaly segmentation by detecting which (if any) of the elements of the sample are anomalous, e.g., by density estimation (Cohen & Hoshen, 2020; Defard et al., 2021; Roth et al., 2022). Given an anomaly segmentation map, we compute the sample-wise anomaly score as the number of anomalous elements, or the abnormality level of the most anomalous element. If the anomaly score exceeds a threshold, the entire sample is denoted as an anomaly. We denote this paradigm detection-by-segmentation.
@@
Existing approaches tackle logical anomalies in images by reconstruction-based approaches - e.g., an autoencoder combining local and global classes (Bergmann et al., 2022; Batzner et al., 2023). These approaches have obtained strong results on some object types, while anomaly detection on other anomaly types remains low. Other approaches provide strong results on all object types but rely on additional supervision (Kim et al., 2024). For the analogues time-series task, where there is a global anomaly in the time series, the best-performing approach is currently a generalization-based approach (Qiu et al., 2021). We aim here to provide strong results on both tasks using a unified density estimation approach.
We propose to detect anomalies consisting of unusual combinations of normal elements using set representations. Our key insight, that we should treat a sample as the set of its elements, is driven by the assumption that in many cases the abnormality of a sample is more correlated with the distribution of elements than with their ordering. Each sample is therefore modeled as an orderless set of elements. The elements are represented using standard fixed feature embeddings, e.g., a deep representation extracted by a pre-trained neural network or handcrafted features. To describe this set of features we model their distribution as a set using a collection of histograms. We compute the histograms for a collection of random projection directions in feature space. The bin occupancies from all the histograms are concatenated together, forming our set representation. Finally, we score anomalies using density estimation on this set representation. We compare our set descriptor to previous approaches and highlight its connection to the sliced Wasserstein distance (SWD).
Our method, SINBAD (Set INspection Based Aomalies Detection) is evaluated on two diverse tasks. The first task is image-level logical anomaly detection on the MVTec-LOCO datasets. We also evaluate our method on series-level time series anomaly detection. Our method outperforms more complex state-of-the-art methods while not using augmentations or training. Note that our method relies on the prior assumption that the elements are normal but their combination is anomalous. In scenarios where the elements themselves are anomalous, it is typically better to perform anomaly detection directly at the element level; E.g., detection-by-segmentation or other methods.
@@ 2 Previous work
Image Anomaly Detection. A comprehensive review of anomaly detection can be found in Ruff et al. (2021). Early approaches such as Glodek et al. (2013); Latecki et al. (2007); Eskin et al. (2002) used handcrafted representations. Deep learning has provided a significant improvement on such benchmarks (Larsson et al., 2016; Ruff et al., 2018; Golan & El-Yaniv, 2018; Hendrycks et al., 2019; Ruff et al., 2019; Perera & Patel, 2019; Salehi et al., 2021; Tack et al., 2020). As density estimation methods utilizing pre-trained deep representation have made significant steps towards the supervised performance on such benchmarks (Deecke et al., 2021; Cohen & Avidan, 2022; Reiss et al., 2021; Reiss & Hoshen, 2021; Reiss et al., 2022); much research is now directed at other challenges (Reiss et al., 2022). Such challenges include detecting anomalous image parts which are small and fine-grained (Cohen & Hoshen, 2020; Li et al., 2021; Defard et al., 2021; Roth et al., 2022; Horwitz & Hoshen, 2022). The progress in anomaly detection and segmentation has been enabled by the introduction of appropriate datasets (Bergmann et al., 2019; 2021; Carrera et al., 2016; Jezek et al., 2021; Bonfiglioli et al., 2022). Recently, the MVTec-LOCO dataset Bergmann et al. (2022) has put the spotlight on fine-grained anomalies that cannot be identified using single patches, but can only be identified when examining the connection between different (otherwise normal) elements in an image. Here, we will focus on detecting such logical anomalies.
@@
Time series Anomaly detection. A general review on anomaly detection in time series can be found in (Blázquez-García et al., 2021). In this paper, we are concerned with anomaly detection of entire sequences, i.e., cases where an entire signal may be abnormal. Traditional approaches for this task include generic anomaly detection approaches such as $k$ nearest neighbors ($k$NN) based methods, e.g., vanilla $k$NN Eskin et al. (2002) and Local Outlier Factor (LOF) Breunig et al. (2000), tree-based methods Liu et al. (2008), one-class classification methods Tax & Duin (2004) and SVDD Schölkopf et al. , and auto-regressive methods that are particular to time series anomaly detection Rousseeuw & Leroy (2005). With the advent of deep learning, the traditional approaches were augmented with deep-learned features: Deep one-class classification methods include DeepSVDD Ruff et al. (2018) and DROCC Goyal et al. (2020). Deep auto-regressive methods include RNN-based prediction and auto-encoding methods Bontemps et al. (2016); Malhotra et al. (2016). In addition, some deep learning anomaly detection approaches are conceptually different from traditional approaches. These methods use classifiers trained on normal data, assuming they will struggle to generalize to anomalous data (Bergman & Hoshen, 2020; Qiu et al., 2021).
@@
Discretized Projections. Discretized projections of multivariate data have been used in many previous works. Locally sensitive hashing Dasgupta et al. (2011) uses random projection and subsequent binary quantization as a hash for high-dimensional data. It was used to facilitate fast $k$ nearest neighbor search. Random projections transformation is also highly related to the Radon transform Radon (1917). Kolouri et al. Kolouri et al. (2015) used this representation as a building block in their set representation. HBOS Goldstein & Dengel (2012) performs anomaly detection by representing each dimension of multivariate data using a histogram of discretized variables. Rocket and mini-rocket Dempster et al. (2020; 2021) represent time series for classification using the averages of their window projection. LODA Pevnỳ (2016) extends this work, by first projecting the data using a random projection matrix. We differ from LODA in the use of a different density estimator and in using sets of multiple elements rather than single sample descriptions.
@@ 3.1 A Set is More Than the Sum of its Parts
Detecting anomalies in complex samples consisting of collections of elements requires understanding how the different elements of each sample interact with one another. As a motivating example let us consider the screw bag class from the MVTec-LOCO dataset (Fig. 1). Each normal sample in this class contains two screws (of different lengths), two nuts, and two washers. Anomalies may occur, for example, when an additional nut replaces one of the washers. Detecting anomalies such as these requires describing all elements within a sample together, since each local element on its own could have come from a normal sample.
A typical way to aggregate element descriptor features is by average pooling - taking the average of the features describing each element. Yet, this is not always suitable for set anomaly detection. In supervised learning, average pooling is often built into architectures such as ResNet He et al. (2016) or DeepSets Zaheer et al. (2017), in order to aggregate local features. Therefore, deep features learned with a supervised loss are already trained to be effective for pooling. However, for lower-level feature descriptors this may not be the case. As demonstrated in Fig.2, the average of a set of features is far from a complete description of the set. This is especially true in anomaly detection, where density estimation approaches require more discriminative features than those needed for supervised learning (Reiss et al., 2022). Even when an average pooled set of features works for a supervised task, it might not work for anomaly detection.
Therefore, we choose to model a set by the distribution of its elements in the embedded feature space, ignoring the ordering between them. A naive way of doing so is using a discretized, volumetric representation, similar to $3$D voxels for point clouds. Unfortunately, such approaches cannot scale to high dimensions, and more compact representations are required. We choose to represent sets using a collection of $1$D histograms. Each histogram represents the density of the elements of the set when projected along a particular direction. We take the bin occupancies of such histograms as our features. We provide an illustration of this idea in Figure 2.
In some cases, projecting a set along its original axes may not be discriminative enough. Histograms along the original axes correspond to 1D marginals, and may map distant elements to the same histogram bins (see Fig.2 for an illustration). On the other side, we can see at the bottom of the figure that when the set elements are first projected along another direction, the histograms of the two sets are distinct. This suggests a set description method: first project each set along a shared random direction and then compute a 1D histogram for each set along this direction. We can obtain a more powerful descriptor by repeating this procedure with projections along multiple random directions. We benchmark this approach in section 3.5.
@@ 3.2 Preliminaries
We are provided a training set $\mathcal{S}$ containing a set of $N_{S}$ samples, we denote a sample as $x\in\mathcal{S}$. We assume that all the training samples are normal. We wish to learn a model that operates on a new, test sample $\tilde{x}$ and outputs an anomaly score. We label samples with anomaly scores higher than a predetermined threshold value as anomalies. The unique aspect of our method is that it treats each sample $x$ as consisting of a set of $N_{E}$ elements, where we denote each element as $e\in x$. Examples of such elements include patches for images, or temporal windows for time series. We assume the existence of a powerful feature extractor $F$ that maps each raw element $e$ into an element feature descriptor $F(e)$. We will describe specific implementations of the feature extraction for two important applications: images and time series, in section 4.
@@ 3.3 Set Features by Histogram of Projections
Motivated by the toy example in section 3.1, we propose to model each set by the histogram of the values of its elements along a collection of directions. We provide an algorithm box Alg.1 summarizing our steps.
Feature extractor. We split each sample to elements $e\in x$ and extract a feature representation for each $\{F(e)\mid e\in x\}$. We describe the implementation of $F$ in Sec.4 as it differs for the time series and image modalities.
Histogram descriptor. While average pooling the features of all elements in the set is an obvious set descriptor, it may result in insufficiently informative representations (section 3.1). Instead, we describe the set by computing the set’s histogram for each feature dimension and concatenating them.
For each sample, we denote the set of values for the $j^{\text{th}}$ dimension of the feature embeddings as $s[j]=\{F(e)[j]\mid e\in x\}$. Note that each set $s[j]$ consists of $N_{E}$ scalar elements coming for each sample. We compute the maximal and minimal values of $s[j]$ for each dimension $j$, among all the elements from all samples combined ($N_{S}\cdot N_{E}$), and divide the region between them into $K$ bins. Finally, we compute histograms $h[j]$ for each of the $N_{D}$ dimensions, describing the set $s[j]$, and concatenate each histogram of $K$ bins of each histogram to a single set descriptor $h\in\mathbb{R}^{N_{D}\cdot K}$.
Projection. As discussed before, not all projection directions are equally informative for describing the distributions of sets. In the general case, it is unknown which directions are the most informative for capturing the difference between normal and anomalous sets. As we cannot tell the best projection directions in advance, we randomly project the features. This minimizes the likelihood of all projections being in catastrophically poor directions, such as those illustrated in Fig. 2.
In practice, we generate a random projection matrix $P\in\mathbb{R}^{N_{D}\times N_{P}}$ by sampling values for each dimension from the Gaussian distribution $\mathbb{N}(0,1)$. We project the features of each element of $x$, yielding projected features $f$:
$f=P\cdot F(e)$ (1)
We run the histogram descriptor procedure described above on the projected features. The final set descriptor $h\in\mathbb{R}^{N_{P}\cdot K}$ is the concatenation of the $N_{P}$ histograms.
@@ 3.5 Connection to Previous Set Descriptors and the Wasserstein Distance
Classical set descriptors. Many prior methods have been used to describe sets of image elements, among them Bag-of-Features Csurka et al. (2004), VLAD Jégou et al. (2010), and Fisher-Vectors Sánchez et al. (2013). These methods begin with a preliminary clustering stage (K-means or Gaussian Mixture Model). They then describe the set using the zeroth, first, or second moments of each cluster. The comparison in Appendix B shows that our method outperforms clustering-based methods in describing our feature sets.
Wasserstein distance. Our method is closely related to the Wasserstein distance, which measures the minimal distance required to transport the probability mass from one distribution to the other. As computing the Wasserstein distance for high-dimensional data such as ours is computationally demanding, the Sliced Wasserstein Distance (SWD) Bonneel et al. (2015), was proposed as an alternative. The $SWD_{1}$ between two sets, $x$ and $y$, has a particularly simple form:
@@
As the histogram projections have a high correlation between them, it is necessary to decorrelate them. This is done here using a Gaussian model. We refer to Appendix G for an explanation of why a Gaussian model is appropriate here. The Mahalanobis distance used here therefore performs better than the simple $SWD_{1}$ distance. While this weakens the connection to the Wasserstein distance, this was crucial for most time-series datasets (see Table 12). In practice, we opted to use $k$NN with the Mahalanobis distance rather than simply computing the Mahalanobis distance to $\mu$ as it worked slightly better (see Appendix B).
We note that the Sliced Wasserstein distance can be calculated directly, without using the suggested histogram binning. However, our histogram-based approach is necessary for our density estimation method. Namely, to compute $k$NN using the Mahalanobis distance, we require a feature representation for each point, rather than relying just on a pairwise distance function between samples. We compare our approach directly with the precise and binned Sliced Wasserstein distance in Appendix F.
Algorithm 1 Set-based Anomaly Detection with Histogram Projections
1:Input: Training set $\mathcal{S}$ with $N_{S}$ normal samples, feature extractor $F$, feature dimension $N_{D}$, number of projections $N_{P}$, number of histogram bins $K$
2:Output: An anomaly detection score
3:Generate random projection matrix $P\in\mathbb{R}^{N_{D}\times N_{P}}$
4:for each sample $x\in\mathcal{S}$ do
5:     for each element $e\in x$ do
6:         Extract features $F(e)$
7:         Project features: $f=P\cdot F(e)$
8:     end for
9:     for each projection dimension $j=1$ to $N_{P}$ do
10:         Build and histogram for each projected dimension, each with $K$ bins
11:         For each sample $x$ we populate the histogram $h_{x}[j]$ of projected features in the $j^{\prime}th$ dimension with the projected features $f[j]$triangleright Each histogram is populated with $N_{E}$ values, as the number of elements extracted from each sample
12:     end for
13:     Concatenate histograms for each sample: $h_{x}=[h_{x}[1],\ldots,h_{x}[N_{P}]]$triangleright resulting dimension is $K\cdot N_{P}$
14:end for
15:Compute mean $\mu$ and covariance $\Sigma$ of $h_{S}$ (the set $h_{S}=\{h_{x}\mid x\in S\}$)
16:for test sample $\tilde{x}$ do
17:     Compute $h_{\tilde{x}}$ as above
18:     Calculate anomaly score: $a(h_{\tilde{x}})=(h_{\tilde{x}}-\mu)^{T}\Sigma^{-1}(h_{\tilde{x}}-\mu)$
19:end for